Change the equation -2x + 8y =4 to slope intercept form.
Here are the steps I have preformed....
8y = 4 + 2y
y = 1/2 + 1/4
Is this right?

Answer by alexander164(3)

All equations are in the form : y=mx+b
ANSWER: -2x+8y=4
8y=2x+4
8y/8=2x/8+4/8
y=1/4x+1/2
Good Luck With Your Homework :)